OpenStudy (anonymous):

1. Solve the following system of equations. -3x + 4y = 80 x + 4y = 32 a. x = 2, y = 21.5 b. x = -12, y = 11 c. x = 11, y = -12 d. x = 8, y = 6

OpenStudy (hoblos):

-3x + 4y = 80 x + 4y = 32 subtract the two equations so you get -4x = 48 x=-12 -3(-12) +4y = 80 4y = 44 y=11